OpenStudy (anonymous):

The equation of line XY is (y−3) = negative 2 over 3(x − 4). What is the slope of a line perpendicular to XY? −2 3 -2/3 3/2 please explain how u got the answer. do not understand.

OpenStudy (anonymous):

do you mean y-3= -2/3 (x-4)? because otherwise your question looks like a rational function to me

OpenStudy (anonymous):

yes lol sorry thats they way they have it worded so its got me confused >_<

OpenStudy (anonymous):

oh ok then its the last one again m= 3/2 because the perpendicular slope of an equation is always the opposite of it.

OpenStudy (anonymous):

and if you're having trouble with these questions I just found a great site that provides an explanation with steps and the solution to similar problems :) in case you still have similar questions http://www.algebra.com/algebra/homework/Linear-equations/equation-parallel-or-perpendicular.solver

OpenStudy (anonymous):

how would i put it into the site?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

like i tried putting the problem into the website u gave me and I'm confused as to what it supposed to be coming out.

OpenStudy (anonymous):

give me a problem ill put it in and link you , if you put the values in the slot provided and scroll all the way down you would see the answer

OpenStudy (anonymous):

y-3= -2/3 (x-4)

OpenStudy (anonymous):

I cant seem to link it but you see if you dont have 2 points to put into the given space then just pick any you want like ex. (0,0) because the slope will stay the same regardless you dont need to worry about that, if youre asked to just find the parallel or perpendicular slope then just pick that option and put in the given slope yu dont need to worry about the point or even the vertical shift for that matter